OpenStudy (anonymous):

Factor completely: 3x2 − 6x − 9

OpenStudy (anonymous):

(3x+3)(x-3)

OpenStudy (anonymous):

x(x+1)(x-3)

OpenStudy (anonymous):

First step, pull the 3 out. So you get 3(x^2-2x-3). Next step is to factor the x^2-2x-3 part. Find two numbers that multiply to -3 and sum to -2.

OpenStudy (anonymous):

- 3x² + 6x - 9 = - 3 ( x² - 2x + 3 ) - 3 x ² = - 3, x, x 6x = 3, 2, x - 9 = - 3, 3 - 3 is the common factor, ( - 3 because there are two negative 3's, and only one positive 3 ) My answer : - 3 ( x² - 2x + 3 )